# xBound: Join Size Lower Bounds

### v1

@@ 3.2. Enabling Reverse Inequalities
Intuition. Let us start with the simplest case, the two-way join. Let $A(X,\ldots)$ and $B(X,\ldots)$ be two relations which join on $X$, and let $D$ be the support of both $A(X)$ and $B(X)$. By the above considerations (§3.1), we know the lower bound $\ell_{A\,\bowtie\,B,\,X,\,0}^{-}\vcentcolon=m$ on the number of distinct $X$ keys of the join $A\bowtie B$. Note that this tells us a little bit more. Indeed, let $\mathbf{a}$ and $\mathbf{b}$ be the degree vectors of column $X$ in $A$ and $B$, respectively. Recall that the lower bound $m$ simply tells us that we can find a set $\mathcal{K}\subseteq A(X)\cap B(X)$ of size $|K|=m$, such that $\mathbf{a}[k]\cdot\mathbf{b}[k]\neq 0,\forall k\in\mathcal{K}$. This, in turn, implies that both $\mathbf{a}$ and $\mathbf{b}$ take positive values on $\mathcal{K}$, which makes them amenable to dot-product reverse inequalities.
@@
Observe now that there is a permutation $\sigma:[|\mathcal{K}|]\to[|\mathcal{K}|]$, such that we can re-write the initial dot-product as
@@
This lemma allows us to connect a lower bound on the number of joining keys with a dot-product on the degree sequences. And this is a key observation: We can now apply any dot-product reverse inequality, as degree sequences have positive entries only.
Example. Let us consider a concrete example. We would like to apply the Pólya–Szegő inequality, Ineq. (2), and we know a lower bound $m=4$ on the number of joining keys. Consider the two degree sequences below:
@@
The lower bound using the degree sequences reads: 1 $\times$ 3 + 1 $\times$ 2 + 1 $\times$ 1 + 2 $\times$ 1 = 8, as we only consider the prefix of length $m=4$. Recall, however, that we only operate with statistics about these degree sequences. Indeed, to apply the Pólya–Szegő inequality, Ineq. (2), we need the values of $\ell_{2},\ell_{\infty}$, and $\ell_{-\infty}$. Calculating them from the (prefixed) degree sequences, we obtain
$\begin{array}[]{c|c|c|c}\text{degree sequence}&\ell_{2}&\ell_{\infty}&\ell_{-\infty}\\ \hline\cr d^{A}[:4]&\sqrt{7}&2&1\\ d^{B}[:4]&\sqrt{15}&3&1\end{array}$
@@
Note that this is slightly worse than the lower bound via the degree sequences directly. This concludes the example for two-way joins.
The topic of actually providing the base table norms will be the topic of  §4. So far, we simply described how to enable dot-product reverse inequalities, by exploiting the lower bound on the number of joining keys calculated in  §3.1. Next, we will generalize the above mechanism to multi-way joins.
Multi-Way Joins. In the sequel, we will be considering a join over the same attribute $X$, i.e., we have $n$ relations $A^{(i)}(X,\ldots)$ and $\mathbf{a}^{(i)}$ the corresponding degree vectors, $i\in[n]$. In addition, denote the domain $D$ as the support of all $A^{(i)}(X)$. Similar to the two-way setting, we will have to first lower bound the number of joining keys that the join produces. This has already been treated in  §3.1. We next show how to improve on this value by using reverse dot-product inequalities.
@@
The result tells us that we can move away from degree vectors, and solely work with the prefixes (of length $m$) of the corresponding degree sequences. Similarly to the two-way case, we will be able to apply a reverse dot-product inequality, namely the generalized reverse Hölder’s theorem, Ineq. (3).
@@ 3.3. Min-Degree Bound
A natural question is whether the above inequalities are exactly what we need to enable lower bounds. On one hand, the above inequalities have the property of being explanatory, i.e., they will return, as LpBound, an expression which can be debugged. On the other hand, they also require the $\ell_{2}$ norm, which is not available by default on the base tables of a system. Driven by a natural quest for simplicity, we found the following bound, which is reminiscent of the max-degree bound in PANDA (Khamis et al., 2017). We term it min-degree.
Essentially, once we have the lower bound on the number of joining keys, $m$, we can fix a relation $R_{i}$, and assume that all other relations join with $\ell_{-\infty}$ keys each. Formally, this reads as follows:
@@ 4. Base Table Statistics
The previous section enabled the application of reverse dot-product inequalities to the setting of degree sequences, and implicitly the norms associated with them. The remaining open question is how to provide these norms to the inequalities. On a high level, given the prefix length $m$, we need to extract the required norms from the degree sequences. In database terminology, we need prefix norms with random access.
@@
An attentive reader would note that simply using the prefix trick does not actually lead to the exact value of the wished-for norms. Surprisingly, we do not need the actual values. Note that we can further lower-bound the RHS of any reverse dot-product inequality by providing hard lower and upper bounds on the requested norms. For instance, to lower bound the RHS side of the Pólya–Szegő inequality, we need (i) lower bounds on $\ell_{2}$, $\ell_{\infty}$, and $\ell_{-\infty}$, and (ii) upper bounds on $\ell_{\infty}$ and $\ell_{-\infty}$. We could infer these lower bounds on the norms by taking a look at the immediate prefixes sizes around $m$, the provided prefixed length. Indeed, for $m$ = 6, we could take as a lower bound for $\ell_{2}$ the value of length 4, while for an upper bound of $\ell_{\infty}$, we could take the value of the prefix length 8, as the degree sequence is sorted in non-decreasing order. While this suffices for correctness, i.e., these are hard lower and upper bounds on the requested norms, there is a way how to further boost their quality. We call this trick norm stitching.
@@ 4.1. Norm Stitching
4.2. Partitioning
Lower bounds on the number of joining keys based only on simple statistics, such as the $\ell_{0}$-value of each column, and the respective min/max values, can be rather loose. One solution is to zoom in on the range value to understand where the joining keys actually occur. To this end, we introduce a lightweight partitioning trick that first splits the value range of each column into a fixed number $B$ of equi-width bins, and collects the same statistics as in  §3.1. Note that partitions for which we collect min/max values can be independent of the actual physical partitions.
Let us outline the benefits this approach. First, collecting $\ell_{0}$-statistics within the $B$ ranges is lightweight (for small $B$’s). Second, this provides us with more fine-grained lower bounds on the number of joining keys, leading to better lower and upper bounds on the norm values (especially when using norm stitching).
@@ 4.3. Predicate Types
At runtime, to fetch the bounds for a range predicate $A\in[x,y]$, we select (i) for the lower bound, the largest bucket that is contained in $[x,y]$, and (ii) for the upper bound, the smallest bucket that contains $[x,y]$.
@@
Predicates $\times$ Partitioning. To integrate with the partitioning trick, one has to extend the support for the above predicates, by performing the very same trick as described in §4.2 for each MCV and histogram bucket, respectively. We stress that fact that we store per partition solely the statistics related to the distinct key count and the min/max values of the key. In particular, we do not store any $\ell_{p}$-norms. It is indeed possible to extend xBound to do so—with presumably better lower bounds, yet this adds another layer of norms, making the computation of statistics computationally intensive.
5. Evaluation
In this section, we want to experimentally understand (1) whether xBound can unearth any lower bound violations in current database systems, and (2) to what extent it can alleviate the original optimizer’s cardinality underestimates.
Table 1. Overview of a system’s cardinality bounds violations across all supported join subexpressions of a benchmark.
Benchmark System Underest’ed ¡ xBound ¿ LpBound
JOBlight [Uncaptioned image] 85.3% 14.9% 8.3%
[Uncaptioned image] 79.3% 6.9% 0.4%
[Uncaptioned image] 53.4% 0.0% 6.0%
JOBjoin [Uncaptioned image] 91.4% 0.0% 5.0%
[Uncaptioned image] 93.1% 2.7% 0.2%
[Uncaptioned image] 90.7% 0.6% 4.2%
STATS-CEB [Uncaptioned image] 90.1% 1.1% 4.0%
[Uncaptioned image] 70.3% 0.3% 3.4%
[Uncaptioned image] 67.0% 0.0% 5.1%
MSSalesJoin [Uncaptioned image] 10.8% 3.9% 89.2%
Table 2. Correcting systems’ cardinality estimates via lower bound clipping: estimate = max(estimate, xBound). We report the median Q-error for each number of joins, stopping at the largest join size for which clipping still yields improvement.
JOBlight
System 1 join 2 joins 3 joins [Uncaptioned image] 2.34 29.45 417.50 xBound-ed [Uncaptioned image] $\Downarrow$ 2.05 $\Downarrow$ 28.30 417.50 [Uncaptioned image] 2.00 1.61 2.09 xBound-ed [Uncaptioned image] $\Downarrow$ 1.76 1.61 2.09
JOBjoin
System 1 join 2 joins 3 joins [Uncaptioned image] 1.01 2.14 3.39 xBound-ed [Uncaptioned image] 1.01 2.14 3.39 [Uncaptioned image] 2.40 2.34 2.25 xBound-ed [Uncaptioned image] $\Downarrow$ 1.00 $\Downarrow$ 2.06 $\Downarrow$ 2.06
STATS-CEB
System 1 join 2 joins 3 joins [Uncaptioned image] 6.72 77.23 518.83 xBound-ed [Uncaptioned image] $\Downarrow$ 6.19 77.23 518.83 [Uncaptioned image] 1.46 2.06 5.07 xBound-ed [Uncaptioned image] 1.46 2.06 5.07
In summary, as we will demonstrate in detail below:
  • 1.
    xBound uncovers lower bound violations in the optimizers of DuckDB 1.4, PostgreSQL 18, and Fabric DW. Depending on the benchmark, the number reaches 15%. These findings may issue a warning to the system in question.
  • 2.
    Via clipping with the lower bound, the xBound’ed systems obtain better one-sided Q-errors, thus showing a first evidence that underestimates could be alleviated before they even happen.
5.1. Experimental Setup
Competitors
We use the traditional optimizers of DuckDB 1.4, PostgreSQL 18, and Fabric Data Warehouse (DW). Notably, these optimizers do not come with any guarantee on the estimates. In addition, we also consider the state-of-the-art pessimistic cardinality estimator LpBound (Zhang et al., 2025), which returns join size upper bounds, thereby serving as a “mirrored” comparator to the lower bounds produced by xBound.
Benchmarks
We use conventional cardinality estimation benchmarks and one production workload: JOBlight (Kipf et al., 2019b), JOBjoin (Zhang et al., 2025), STATS-CEB (Han et al., 2021), and Microsoft’s MSSalesJoin. JOBjoin is a recent variation of JOB (Leis et al., 2015) without table predicates for measuring the quality of pessimistic cardinality estimators. [^3] We consider only the self-join-free queries, as in the same reference. MSSalesJoin is a join-only variation of the analytical production workload from Microsoft’s internal revenue reporting platform MSSales (Alotaibi et al., 2024), using a 10%-sample, in total 500 GB, of the real-world dataset. This is a significantly larger real dataset compared to the other benchmarks, while the queries have similar complexity to JOBjoin.
For all these benchmarks, we work with the SQL subexpressions we (currently) support, namely multi-way joins on the same key. Surprisingly, this is exactly the setup in JOBlight, the benchmark used par excellence by learned estimators (Kipf et al., 2019b). In addition, in the case of JOBjoin, we pushed the limit a bit further: given that the subexpressions tend to have a multi-way join on the same join-key which, in turn, joins with unfiltered dimension tables, e.g., info_type, we directly report the size of the respective multi-way join. For an overview of the number of supported join subexpressions and their maximum number of joins, consider Tab. 3.
Table 3. Overview of benchmarks used in the experiments: (i) total number of benchmark queries, (ii) ratio of supported subexpressions, and (iii) the maximum number of joins in the supported subexpressions.
Benchmark #$Q$’s %supported-subexpr’s max. #joins
JOBlight 70 696 / 696 4
JOBjoin 31 9566 / 9566 16
STATS-CEB 146 1830 / 3235 4
MSSalesJoin 47 508 / 50k 3
System configuration

### v2

@@ 3.3. Enabling Reverse Inequalities
Intuition. Let us start with the simplest case, the two-way join. Let $A(X,\ldots)$ and $B(X,\ldots)$ be two relations which join on $X$, and let $D$ be the support of both $A(X)$ and $B(X)$. By the above considerations (§3.1), we know the lower bound $\ell_{A\,\bowtie\,B,\,X,\,0}^{-}\vcentcolon=m$ on the number of distinct $X$ keys of the join $A\bowtie B$. Note that this tells us a little bit more. Indeed, let $\mathbf{a}$ and $\mathbf{b}$ be the degree vectors of column $X$ in $A$ and $B$, respectively. Recall that the lower bound $m$ simply tells us that we can find a set $\mathcal{K}\subseteq A(X)\cap B(X)$ of size $|K|=m$, such that $\mathbf{a}[k]\cdot\mathbf{b}[k]\neq 0,\forall k\in\mathcal{K}$. This, in turn, implies that both $\mathbf{a}$ and $\mathbf{b}$ take positive values on $\mathcal{K}$, which makes them amenable to inner-product reverse inequalities.
@@
Observe now that there is a permutation $\sigma:[|\mathcal{K}|]\to[|\mathcal{K}|]$, such that we can re-write the initial inner product as
@@
This lemma allows us to connect a lower bound on the number of distinct joining keys with a inner product on the degree sequences. And this is a key observation: We can now apply any inner-product reverse inequality, as degree sequences have positive entries only.
Example 1. Let us consider a concrete example. We take the simplest reverse inequality, Ineq. (2), which requires only the $\ell_{1}$ and the $\ell_{-\infty}$ statistics. Assume we already computed a lower bound $m=4$ on the number of distinct joining keys. Consider the two degree sequences:
@@
The lower bound using the degree sequences reads: 1 $\times$ 3 + 1 $\times$ 2 + 1 $\times$ 1 + 2 $\times$ 1 = 8, as we only consider the prefix of length $m=4$. Recall, however, that we operate only with the statistics of these degree sequences. Calculating them from the (prefixed) degree sequences, we obtain
$\begin{array}[]{c|c|c}\text{degree sequence}&\ell_{1}&\ell_{-\infty}\\ \hline\cr d^{A}[:4]&{\color[rgb]{0,0,1}1}+{\color[rgb]{0,0,1}1}+{\color[rgb]{0,0,1}1}+{\color[rgb]{0,0,1}2}={\color[rgb]{0,0,1}5}&{\color[rgb]{0,0,1}1}\\ d^{B}[:4]&{\color[rgb]{1,.5,0}1}+{\color[rgb]{1,.5,0}1}+{\color[rgb]{1,.5,0}2}+{\color[rgb]{1,.5,0}3}={\color[rgb]{0,0,1}7}&{\color[rgb]{1,.5,0}1}\end{array}$
Therefore, this tells us that (i) 5 keys from $A$ can join with 1 key from $B$ and (ii) 7 keys from $B$ can join with 1 key from $A$. Taking the maximum of these two options yields a lower bound 7. Note that this is slightly worse than the lower bound via the degree sequences directly.
This is what we refer to as the min-degree bound, and is formalized in the upcoming §3.4. Next, we consider another example, where we consider an $\ell_{2}$-based reverse inequality.
Example 2. We would like to apply the Pólya–Szegő inequality, Ineq. (3), and we know a lower bound $m=4$ on the number of distinct joining keys, as in the previous example. Moreover, the two degree sequences stay the same.
To apply the Pólya–Szegő inequality, Ineq. (3), we need the values of $\ell_{2},\ell_{\infty}$, and $\ell_{-\infty}$. Calculating them from the (prefixed) degree sequences, we obtain
$\begin{array}[]{c|c|c|c}\text{degree sequence}&\ell_{2}&\ell_{\infty}&\ell_{-\infty}\\ \hline\cr d^{A}[:4]&\sqrt{{\color[rgb]{0,0,1}1}^{2}+{\color[rgb]{0,0,1}1}^{2}+{\color[rgb]{0,0,1}1}^{2}+{\color[rgb]{0,0,1}2}^{2}}={\color[rgb]{0,0,1}\sqrt{7}}&{\color[rgb]{0,0,1}2}&{\color[rgb]{0,0,1}1}\\ d^{B}[:4]&\sqrt{{\color[rgb]{1,.5,0}1}^{2}+{\color[rgb]{1,.5,0}1}^{2}+{\color[rgb]{1,.5,0}2}^{2}+{\color[rgb]{1,.5,0}3}^{2}}={\color[rgb]{1,.5,0}\sqrt{15}}&{\color[rgb]{1,.5,0}3}&{\color[rgb]{1,.5,0}1}\end{array}$
@@
This is (slightly) better than the lower bound achieved by the min-degree bound. This concludes the examples for two-way joins.
The topic of actually providing the base table norms will be the topic of  §4. So far, we simply described how to enable inner-product reverse inequalities, by exploiting the lower bound on the number of distinct joining keys calculated in  §3.1. Next, we will generalize the above mechanism to multi-way joins.
Multi-Way Joins. In the sequel, we will be considering a join over the same attribute $X$, i.e., we have $n$ relations $A^{(i)}(X,\ldots)$ and $\mathbf{a}^{(i)}$ the corresponding degree vectors, $i\in[n]$. In addition, denote the domain $D$ as the support of all $A^{(i)}(X)$. Similar to the two-way setting, we will have to first lower bound the number of distinct joining keys that the join produces. This has already been treated in  §3.1. We next show how to improve on this value by using reverse inner-product inequalities.
@@
The result tells us that we can move away from degree vectors and instead work solely with the prefixes (of length $m$) of the corresponding degree sequences. This enables the application of all reverse inner-product inequalities described in the preliminaries, in particular the generalized reverse Hölder’s inequality, Ineq. (4).
@@ 3.4. Min-Degree Bound
As pointed out in our first example, we can use Ineq. (2) to have what we term the min-degree bound, which is reminiscent of the max-degree bound in PANDA (Khamis et al., 2017). Compared to the other reverse inequalities, it does not require the $\ell_{2}$ norm, which is not available by default on database catalogs.
The key idea is that, once we have the lower bound on the number of distinct joining keys, $m$, we can fix a relation $R_{i}$, and assume that all other relations join with $\ell_{-\infty}$ keys each. Formally, this reads as follows:
@@
3.5. Heavy Partition
For a two-way join, frameworks for upper bounds on query sizes, such as LpBound and SafeBound, tend to ignore which keys join from the two tables. And, indeed, this is a pragmatic approach, since keeping track of the joining keys themselves is expensive. In our experiments, however, we saw that lower bounds cannot be competitive enough if we do not maintain, at least, a small set of heavy hitters that are going to be joined. Hence, apart from the (light) partitions we will be introducing later (§4.2), we will maintain on each table a set of heavy hitters, which we refer to as the heavy partition. For a given join pool (recall §2), we maintain a set $\mathcal{H}$ that represents the heavy hitters across the corresponding join columns. We keep this set constant across all supported predicate types, and limit its cardinality to $\mathcal{H}$ to a small constant (set in our experiments to 64). To this end, we used the FrequentItems sketch (Misra and Gries, 1982) available in Apache’s DataSketches library, with a precision of 12, and extracted the heavy keys without false positives, along with a (hard) lower bound on their degrees; this guarantees a single-pass computation.
At estimation time, the heavy partition is tracked of separately, namely: we track which heavy keys survive the join, and then collect their exact degrees (or lower bounds thereof) from the pre-computed statistics. We discuss how to compute these statistics in §4.4. For non-selective queries, integrating a heavy partition into xBound enabled almost perfect estimates (only some constant factors away), while preserving the guarantee of hard lower bounds.
Note that Fabric DW already provides a way to keep track of a column’s heavy hitters and their (approximate) degrees, by employing a CountMinSketch (Cormode and Muthukrishnan, 2005). Hence, xBound’s heavy partition can be supported without much extra overhead.
@@ 4. Base Table Statistics
The previous section enabled the application of reverse inner-product inequalities to the setting of degree sequences, and implicitly the norms associated with them. The remaining open question is how to provide these norms to the inequalities. On a high level, given the prefix length $m$, we need to extract the required norms from the degree sequences. In database terminology, we need prefix norms with random access.
@@
An attentive reader would note that simply using the prefix trick does not actually lead to the exact value of the wished-for norms. Surprisingly, we do not need the actual values. Note that we can further lower-bound the RHS of any reverse inner-product inequality by providing hard lower and upper bounds on the requested norms. For instance, to lower bound the RHS side of the Pólya–Szegő inequality, we need (i) lower bounds on $\ell_{2}$, $\ell_{\infty}$, and $\ell_{-\infty}$, and (ii) upper bounds on $\ell_{\infty}$ and $\ell_{-\infty}$. We could infer these lower bounds on the norms by taking a look at the immediate prefixes sizes around $m$, the provided prefixed length. Indeed, for $m$ = 6, we could take as a lower bound for $\ell_{2}$ the value of length 4, while for an upper bound of $\ell_{\infty}$, we could take the value of the prefix length 8, as the degree sequence is sorted in non-decreasing order. While this suffices for correctness, i.e., these are hard lower and upper bounds on the requested norms, there is a way how to further boost their quality. We call this trick norm stitching.
@@ 4.1. Norm Stitching
4.2. Light Partitions
Lower bounds on the number of joining keys based only on simple statistics, such as the $\ell_{0}$-value of each column, and the respective min/max values, can be rather loose. This holds also for the extension to probabilistic $\ell_{0}$ in lower bounds outlined in §3.2. To this end, we introduced in §3.5 the idea of a heavy partition, which is targeted at heavy hitters. For the rest of the keys, one solution is to zoom in on the range value to understand where the joining keys actually occur. Hence, we introduce a lightweight partitioning trick that first splits the value range of each column into a fixed number $B$ of equi-width bins, and collects the same statistics as in  §3.1. We refer to these as the light partitions. Note that the partitions for which we collect min/max values can be independent of the actual physical table partitions. To exemplify this, note that the light partitions can also be done a via an arbitrary hash function when using the probabilistic lower bounds; the hard lower bounds do need the min/max statistics, as explained §3.1.
Let us outline the benefits this approach. First, collecting $\ell_{0}$ statistics within the $B$ ranges is lightweight (for small $B$’s). Second, this provides us with more fine-grained lower bounds on the number of joining keys, leading to better lower and upper bounds on the norm values that are plugged into the inner-product inequalities, especially when using norm stitching (§4.1).
@@ 4.3. Predicate Types
At runtime, to fetch the bounds for a range predicate $A\in[x,y]$, we select (i) for the lower bound, the largest bucket that is contained in $[x,y]$, and (ii) for the upper bound, the smallest bucket that contains $[x,y]$. However, we observed that, on large datasets, supporting the prefixed norms in this case adds a considerable overhead when pre-computing the statistics. This is due to the fact that we did not exploit the overlaps naturally found in the range buckets. As a tradeoff, we chose the take at estimation time, as a valid lower bound, solely the $\ell_{0}$ lower bound on the corresponding bucket, without relying on the $\ell_{p}$ norms.
@@
Predicates $\times$ Partitioning. To integrate with the partitioning trick, one has to extend the support for the above predicates, by performing the very same trick as described in §4.2 for each MCV and histogram bucket, respectively. We stress that fact that we store per partition solely the statistics related to the distinct key count (or the Theta sketches, in the case of the probabilistic $\ell_{0}$ lower bounds) and the min/max values of the key.
Note that we do not store any $\ell_{p}$-norms. It is indeed possible to extend xBound to do so—with presumably better lower bounds, yet this adds another layer of norms, making the computation of statistics computationally intensive. Since
4.4. Predicates $\times$ Heavy Keys
As a booster for our lower bounds, we introduced in §3.5 the concept of a heavy partition: heavy hitters appearing on the join columns. In the following, we outline how they are supported in combination with the predicates. The statistics are stored the hh_stats table. First, in the case of no predicates, we store per heavy join key its degree (frequency) on each table. Equality predicates are supported via the MCV support: For each MCV on the predicate column, we store in hh_stats the degrees of the heavy keys; similarly for the range predicates, where we store the degree w.r.t. to a range bucket.
Multiple Predicates. While this yields exact lower bounds when having table scans with at most one predicate, the case of multiple predicates incurs a challenge. Namely, we have to lower bound the degree of a heavy key under two or more predicates, which, without any additional information, can only be set to 0. While this can be well pre-computed for two predicates, e.g., storing the degrees under an MCV and a range bucket, this can be quite of an overhead. Instead, we opted for a probabilistic approach: per predicate type, we can store a ThetaSketch on the tuple-id in hh_info, and, at estimation time, lower-bound, with a given confidence, the intersection size of the sets represented by the sketches. Indeed, this gives a probabilistic lower bound on the number of tuples that match all predicates, i.e., on the degree of the respective heavy key. Notably, applying this trick for all heavy keys weakens the guarantees of the lower bounds. We plan to investigate the resulting confidence of these probabilistic lower bounds.
Figure 6. Reducing Fabric DW’s query underestimates via clipping with xBound’s lower bounds on the SO-CEB benchmark. We show the queries in order of their Fabric DW’s Q-error and highlight the queries with improvements.
5. Experiments
We now empirically evaluate the effectiveness of xBound in reducing cardinality underestimation in a real production system, Fabric DW. Particularly, we demonstrate that clipping Fabric DW’s estimates with xBound’s lower bounds directly improves resource estimation, yielding sizable query speedups.
Benchmark. Since Fabric DW is a cloud-based distributed data warehouse, we need a dataset large enough to exercise its full capability. We therefore run the STATS-CEB benchmark (Han et al., 2021) on the 220 GB StackOverflow dump (Schmidt et al., 2025), ensuring distributed operators appear in our system’s query plans; we refer to this as StackOverflow-CEB (SO-CEB). Note that we remove filters on the posts.favoritecount column, as it is NULL in this dump. The benchmark comprises 146 Berge-acyclic queries over up to 7 tables with equality and range predicates on numeric and date columns. [^5] We consider their simplest form: select count(*) from [...]. Of these, xBound currently supports 89 / 146 queries whose (mostly FK-FK) joins share a single join key; specifically, they join either on userid or postid, including semantically equivalent attributes such as posts.owneruserid and postlinks.relatedpostid.